OpenStudy (anonymous):

Would someone explain how to set this problem up? Solve the equation: log(8)x + log(11)x = 1

OpenStudy (saifoo.khan):

8 and 11 are in base?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\Large \log_8 x + \log_{11} x =1\] \[\frac{\log x}{\log 8} + \frac {\log x}{\log 11} = 1\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so then what steps would allow me to solve for x ?

OpenStudy (saifoo.khan):

take lcm.

OpenStudy (anonymous):

88 ?

OpenStudy (saifoo.khan):

\[\frac{ (\log x * \log11) + (\log x * \log 8)}{\log8 * \log11}=1\]

OpenStudy (anonymous):

ok, ty!

OpenStudy (anonymous):

so then x= 8, 11

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

\[\frac {\log x (\log 11 + \log8)}{\log 88} = 1\]

OpenStudy (anonymous):

ok, my problem asks: x= or there is no solution.

OpenStudy (saifoo.khan):

Solve for x. \[\log x = \frac{\log 88}{\log11 + \log8}\]

OpenStudy (saifoo.khan):

Insert the values in a calc.

OpenStudy (anonymous):

2.770